# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v2

@@ Abstract
Our thorough and multifaceted analysis shows that, in controlled, toy settings with synthetic data, our approach can recover the ground-truth encoding-decoding direction pairs. In real-world settings, our method effectively reveals the encoding-decoding mechanism of interpretable concepts, often scoring substantially better in interpretability metrics than other unsupervised baselines. Finally, we provide concrete applications of how the learned directions can help open the black box and understand global model behavior, explain individual sample predictions in terms of local, spatially-aware, concept contributions and intervene on the network’s prediction strategy to provide either counterfactual explanations or correct erroneous model behavior.
@@ 1 Introduction
The linear representation hypothesis suggests that deep neural networks encode high-level concepts in directions of their latent space (mi-review). This hypothesis is sufficiently supported by empirical evidence, mostly by the effectiveness of linear probing. In the latter, a single linear layer can be trained on top of feature representations originating from the upper part of deep neural networks to solve semantic tasks with great success (IntriguingProperties; ClassifierProbes; IBD; TCAV; Superposition; LinearRepresentationsNLP). In the concept encoding mechanism suggested by the linear representation hypothesis, the concept content of an input patch is written to its embedding as a linear combination of concept embeddings: directions in the latent space represented as vectors, each one associated with a concept. The corresponding scalar coefficient in the linear combination constitutes the concept’s latent factor that encodes the degree of concept presence in the input. Due to its utility in encoding the latent factor to a vector representation, we also refer to the concept embedding as the concept’s encoding or signal direction. Retrieving the latent factor back can be accomplished by taking the inner product between the patch embedding and a filter, another vector which we term a decoding direction. (Fig. 1) An encoding-decoding direction pair is interpretable whenever it is related to a concept that is aligned with human intuition; and is influential whenever it is related to a concept that the network consistently uses to make predictions (TCAV; RCAV).
The encoding - decoding mechanism of concepts for a pre-trained network is not directly accessible. Instead, it is a latent mechanism emerged from network training and needs to be inferred by any means of reverse engineering. Recovering this mechanism is of significant importance, as accessing it enables understanding representations (IBD; TCAV), assessing influence of concepts on the network’s predictions (TCAV; RCAV; CRAFT; PCAV), debugging neural networks (XAIImageNet), and correcting erroneous model behavior (RemovingCleverHans; RevealToRevise; spurious-features; GradientPenalization).
Previous works that identify concept directions rely either on concept annotations (IBD; TCAV; PCAV) or on a decompose-then-reconstruct regime of feature activations (NMFCAV; SVD; DictionaryNLP; SAE1; SAE2; matryoshka-saes). The former suffer from the need for concept speculation and the access to annotations while the latter, face the difficulty of overcoming unstructured noise that may be present in feature activations in the form of distractors (NeuroImaging; PatternNet) or dark matter (SAE-dark-matter).
@@
Unlike previous approaches, our method relies on neither annotations nor the reconstruction of activations, thereby sidestepping these earlier limitations. Instead, we identify decoding directions via directional clustering of feature activations and estimate encoding directions from a probabilistic perspective. Moreover, in contrast to most prior work, we guide our search for directions by exploiting the network’s instructions, i.e. the prediction strategy baked into its weights through a process we call Uncertainty Region Alignment. We conduct a thorough and multifaceted analysis that highlights the effectiveness of the approach in revealing the encoding-decoding mechanism of a) the ground-truth concepts in controlled and synthetic toy settings and b) interpretable monosemantic concepts in real-world studies of state-of-the-art deep vision architectures We also demonstrate the utility of our learned directions in obtaining global, local, and counter-factual explanations, as well as identifying and correcting flawed model behavior in a toy setting (Fig. 2).
@@ 2.1 Direction Learning
Concept Activation Vectors (CAVs) (TCAV; IBD), use supervised linear classifiers trained on annotated concept datasets to separate representations of samples with and without a target concept. To solve the task, the resulting classifier weights (filters) must extract (decode) the concept factors from these representations; thus, they approximate concept decoding directions, although they can be inexact due to distractor-noise in the features (NeuroImaging; PatternNet; PCAV). To estimate concept encoding directions, PCAVs (PCAV) take the difference of means between representations of samples with and without the concept. These methods require both labeled data and prior specification of the concept name. In contrast to these probing approaches, the proposed method is unsupervised, revealing concepts by analyzing the structure of the latent space, and thereby avoiding both annotation costs and upfront concept specification.
@@
Unsupervised Concept Direction Learning Existing unsupervised direction-learning methods that rely on matrix decomposition (PCA (PCA), SVD (SVD), NMF (NMFCAV; CRAFT)) can discover concept directions without annotations but have key limitations. PCA and SVD are constrained by orthogonality and miss concepts that do not influence variance (ConceptExtractionAndImportance). NMF, while non-negative, lacks latent space bias and, thus, expressivity, and it has no straightforward decoding directions. The proposed method addresses and overcomes all these limitations.
@@
Dictionary Learning (DictionaryNLP; DictionaryNLP2) and Sparse Autoencoders (SAEs) (SAE3; SAE2; matryoshka-saes) identify encoding-decoding directions driven by the reconstruction of representations after decomposing them into sparse latent factors. They typically consider a latent space larger than the embedding space and may suffer from the difficulty of overcoming noise in the representations in the form of dark matter (SAE-dark-matter). Contrariwise, our method is independent of feature reconstruction, unrestricted from the relative sizes between the concept and the embedding space, and unlike these prior works, it implicitly links the learned directions to how the model actually utilizes them. Further comparative details with SAEs are provided in Section A.22.
@@
Moving away from reconstructing representations, the methods of UIBE; CBE perform directional clustering of feature activations, a process that often unveils the decoding mechanism of interpretable, monosemantic concepts. We build on these previous approaches to estimate decoding directions and do so by removing orthogonality and feature standardization constraints, and adding new loss terms to maintain or improve concept interpretability and reduce the impact of distractor noise in filter weights. Compared to the previous attempt (CBE) of linking directions to their utilization by the network, our Uncertainty Region Alignment approach achieves up to 22.56% relative improvement in interpretability metrics in 3 out of 4 cases. We also address the estimation of concept encoding directions, which these prior works did not, with further details on this comparison provided in Section A.3.
@@ 2.2 Applications of Directions
Our learned direction pairs enable a series of applications, from global model understanding and detailed spatially-aware local explanations down to model intervention, which allows for counterfactual explanations and model correction. Typical previous enablers of such applications are the Concept Bottleneck Models (CBMs) (CBMs). Classic CBMs consist of a backbone feature extractor, a concept bottleneck predicting concept content from features, and a linear head mapping concepts to classes. What makes this architecture interpretable is the fact that the concept vector, i.e, the output of the concept bottleneck, is trained to have each dimension aligned with an interpretable concept. Recent post-hoc CBM variants (PostHocCBM; LabelFreeCBM) retrofit this structure onto existing black-box models by freezing the original backbone, discarding its head, inserting a concept bottleneck projection into concept space, and learning a task-specific sparse linear head over concepts. To learn concepts without annotations, these works exploit CLIP CLIP, and further advances spatial-cbms use visual prompting (clip-visual-prompt) to make the bottleneck spatially aware, allowing localization of concepts at the patch level rather than only at the image level. Most recently (discover-then-name-sae-cbms), SAEs (SAE2) have been used to build CBMs directly for CLIP.
To access the advantages of CBMs, these previous approaches require training the last linear head using the model’s training data and, sometimes, additional concept-annotated datasets or external models like CLIP. In contrast, the proposed work is non-intrusive: it needs no training of new model components and no concept labels. It shows that one can capture many benefits of concept bottleneck models by identifying directions corresponding to concepts the model already encodes and then using these directions in applications of mechanistic interpretability (mechanistic; mechanistic-interpretability-1; mi-review).
@@ 3 Background
In our formulation, which follows the linear representation hypothesis (Superposition; LinearRepresentationsNLP), the latent factor of a concept is a scalar denoting the concept’s presence in the input, and is embedded in the latent space via multiplication with the concept’s encoding direction, also called the concept’s embedding or signal direction (PatternNet). For this reason, we also refer to this latent factor as the signal value. Feature activations are considered as linear combinations of signals and noisy directional components called distractors. In the proposed approach, a filter, through the inner product with an input representation, can extract the signal’s value and for this reason we term it a decoding direction. Below we provide a more formal explanation of these terms and provide details essential to understand our contributions.
@@ 3.2 Signals, Distractors, Filters, Concept-Detectors
In encoding a single concept $i$, PatternNet; PCAV proposed a model for the data generation process of feature representations: ${\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{d}},\,{\bm{s}}_{i},{\bm{d}}\in{\mathbb{R}}^{D},\alpha_{{\bm{p}}},\beta_{{\bm{p}}}\in{\mathbb{R}}$. Here, ${{\bm{s}}}_{i}$ is the signal direction that carries the information of whether ${\bm{x}}_{{\bm{p}}}$ is part of concept $i$. The concept information lies within the signal value $\alpha_{{\bm{p}}}$. Larger $\alpha_{{\bm{p}}}$ suggests greater confidence that ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. $\bm{d}$} is the distractor direction, modeling noise, or information not related to the concept. $\beta_{{\bm{p}}}$ follows a random distribution, typically the gaussian or uniform distribution, and is independent of whether ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. According to PatternNet, the value of the signal $\alpha_{{\bm{p}}}$ can be extracted using a filter ${\bm{w}}_{i}$ and the inner product: $z_{{\bm{p}},i}={\bm{w}}_{i}^{T}{\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{\bm{w}}_{i}^{T}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{w}}_{i}^{T}{\bm{d}}$, if we choose ${\bm{w}}_{i}$ : ${\bm{w}}_{i}\perp{\bm{d}}$, and ${\bm{w}}_{i}^{T}{{\bm{s}}}_{i}=1$. Let $\sigma$ denote the sigmoid function. Since stronger values of $\alpha_{{\bm{p}}}$ indicate more confidence in concept presence, when combined with a threshold $b_{i}\in{\mathbb{R}}$ that can be learned from data, this filter can be turned into a concept detector: $y_{{\bm{p}},i}=\sigma(z_{{\bm{p}},i}-b_{i})$, essentially a binary classifier that can answer the question of whether $\bm{p}$} belongs to concept $i$.
With access to the signal value, NeuroImaging; PatternNet offer a formula to estimate the concept’s signal direction:
@@ 3.3 Unsupervised Interpretable Direction Learning
Figure 3: The core concept of Unsupervised Interpretable Basis Extraction (UIBE UIBE) is to learn a set of concept detectors, which are essentially binary linear classifiers with learnable filters and biases. These detectors aim to transform feature representations to the soft-binary vector space of concepts in which the newly transformed representations are sparse. In this procedure the input to the method consists of image representations coming from an unlabeled concept dataset. Identifying the concept name behind each detector is done in a post-processing step with a procedure we refer to as Direction Labeling.
Recent research (UIBE) introduced an unsupervised method to identify concepts from the structure of the latent space. Motivated by the directional encoding of concepts, the method partitions the latent space into linear regions, each represented by a hyperplane and a normal vector, forming clusters. Feature activations from an unlabeled concept dataset, possibly activations of images from the network’s training set, are assigned to these clusters. The method learns $\bm{W}$} and $\bm{b}$} of a feature-to-cluster membership function, a mapping to the semantic space, with ${\bm{y}}_{{\bm{p}}}=\sigma({\bm{W}}^{T}{\bm{x}}_{{\bm{p}}}-{\bm{b}})\in[0,1]^{I},{\bm{W}}\in{\mathbb{R}}^{D\times I},{\bm{b}}\in{\mathbb{R}}^{I}$, and $I$ as the cluster count. By softly assigning features to a small number of clusters, the interpretability of the clustering is improved. This is grounded in the idea that an image patch generally holds only a few semantic labels from a larger set, reflecting sparsity in the semantic space. Sparsity in the assignments is achieved using two loss terms: the first is Sparsity Loss ($\mathcal{L}^{s}$), and the second is Maximum Activation Loss ($\mathcal{L}^{ma}$), which ensures binary cluster membership:

### v3

@@ Abstract
Our thorough and multifaceted analysis shows that, in controlled, toy settings with synthetic data, our approach can recover the ground-truth encoding-decoding direction pairs. In real-world settings, our method effectively reveals the encoding-decoding mechanism of interpretable concepts, often scoring substantially better in interpretability metrics than other unsupervised baselines, such as PCA and NMF. Finally, we provide concrete applications of how the learned directions can help open the black box and understand global model behavior, explain individual sample predictions in terms of local, spatially-aware, concept contributions and intervene on the network’s prediction strategy to provide either counterfactual explanations or correct erroneous model behavior.
@@ 1 Introduction
The linear representation hypothesis suggests that deep neural networks encode high-level concepts in directions of their latent space (Bereska and Gavves, 2024). This hypothesis is sufficiently supported by empirical evidence, mostly by the effectiveness of linear probing. In the latter, a single linear layer can be trained on top of feature representations originating from the upper part of deep neural networks to solve semantic tasks with great success (Szegedy et al., 2014; Alain and Bengio, 2018; Zhou et al., 2018; Kim et al., 2018; Elhage et al., 2022; Nanda et al., 2023). In the concept encoding mechanism suggested by the linear representation hypothesis, the concept content of an input patch is written to its embedding as a linear combination of concept embeddings: directions in the latent space represented as vectors, each one associated with a concept. The corresponding scalar coefficient in the linear combination constitutes the concept’s latent factor that encodes the degree of concept presence in the input. Due to its utility in encoding the latent factor to a vector representation, we also refer to the concept embedding as the concept’s encoding or signal direction. Retrieving the latent factor back can be accomplished by taking the inner product between the patch embedding and a filter, another vector which we term a decoding direction. (Fig. 1) An encoding-decoding direction pair is interpretable whenever it is related to a concept that is aligned with human intuition; and is influential whenever it is related to a concept that the network consistently uses to make predictions (Kim et al., 2018; Pfau et al., 2020).
The encoding - decoding mechanism of concepts for a pre-trained network is not directly accessible. Instead, it is a latent mechanism emerged from network training and needs to be inferred by any means of reverse engineering. Recovering this mechanism is of significant importance, as accessing it enables understanding representations (Zhou et al., 2018; Kim et al., 2018), assessing influence of concepts on the network’s predictions (Kim et al., 2018; Pfau et al., 2020; Fel et al., 2023b; Pahde et al., 2024), debugging neural networks (Anders et al., 2020), and correcting erroneous model behavior (Anders et al., 2022; Pahde et al., 2023; Neuhaus et al., 2023; Dreyer et al., 2024a).
Previous works that identify concept directions rely either on concept annotations (Zhou et al., 2018; Kim et al., 2018; Pahde et al., 2024) or on a decompose-then-reconstruct regime of feature activations (Zhang et al., 2021; Graziani et al., 2023b; Bricken et al., 2023; Lim et al., 2024; Cunningham et al., 2024; Bussmann et al., 2025). The former suffer from the need for concept speculation and the access to annotations while the latter, face the difficulty of overcoming unstructured noise that may be present in feature activations in the form of distractors (Haufe et al., 2014; Kindermans et al., 2017) or dark matter (Engels et al., 2025).
@@
Unlike previous approaches, our method relies on neither annotations nor the reconstruction of activations, thereby sidestepping these earlier limitations. Instead, we identify decoding directions via directional clustering of feature activations and estimate encoding directions from a probabilistic perspective. Moreover, in contrast to most prior work, we guide our search for directions by exploiting the network’s instructions, i.e. the prediction strategy baked into its weights through a process we call Uncertainty Region Alignment. We conduct a thorough and multifaceted analysis that highlights the effectiveness of the approach in revealing the encoding-decoding mechanism of (a) the ground-truth concepts in controlled and synthetic toy settings, and (b) interpretable monosemantic concepts in real-world studies of state-of-the-art deep vision architectures. We also demonstrate the utility of our learned directions in obtaining global, local, and counter-factual explanations, as well as identifying and correcting flawed model behavior in a toy setting (Fig. 2).
Our comparative analysis primarily considers established matrix decomposition baselines, specifically Principal Component Analysis (PCA) and Non-Negative Matrix Factorization (NMF). We further extend the comparisons in Appendix Section A.16, where we evaluate our approach against the methods on which we build, namely Unsupervised Interpretable Basis Extraction (UIBE) (Doumanoglou et al., 2023) and Concept Basis Extraction (CBE) (Doumanoglou et al., 2024). Throughout, we focus on settings in which the concept space does not exceed the dimensionality of the embedding space, and we leave a systematic comparison with Sparse AutoEncoders (SAEs) (Sharkey et al., 2022) to future work.
@@ 2.1 Direction Learning
Concept Activation Vectors (CAVs) (Kim et al., 2018; Zhou et al., 2018) use supervised linear classifiers trained on annotated concept datasets to separate representations of samples with a target concept from those without. To solve the task, the resulting classifier weights (filters) must extract (decode) the concept factors from these representations; thus, they approximate concept decoding directions, although they can be inexact due to distractor-noise in the features (Haufe et al., 2014; Kindermans et al., 2017; Pahde et al., 2024). To estimate concept encoding directions, PCAVs (Pahde et al., 2024) take the difference of means between representations of samples with and without the concept. These methods require both labeled data and prior specification of the concept name. In contrast to these probing approaches, the proposed method is unsupervised, revealing concepts by analyzing the structure of the latent space and thereby avoiding both annotation costs and upfront concept specification.
@@
Unsupervised Concept Direction Learning Existing unsupervised direction-learning methods that rely on matrix decomposition (PCA (Graziani et al., 2023a), SVD (Graziani et al., 2023b), NMF (Zhang et al., 2021; Fel et al., 2023b)) can discover concept directions without annotations, but they have key limitations. PCA and SVD are constrained by orthogonality and miss concepts that do not influence variance (Fel et al., 2023a). NMF, while non-negative, lacks latent space bias and, thus, expressivity, and it has no straightforward decoding directions. To alleviate these shortcomings, at least to some extent, Graziani et al. (2023a; b) combined direction learning with gradient re-ranking, and both Zhang et al. (2021) and Graziani et al. (2023a; b) worked on subgroups of features belonging to the same class category, establishing a class-conditioned protocol for direction learning. In contrast, the proposed method uses a class-agnostic protocol without the need for sample class labels. Additionally, our approach considers non-orthogonal concept directions and accounts for feature space bias, overcoming the limitations of previous methods.
@@
Dictionary Learning (Bricken et al., 2023; Yun et al., 2023) and Sparse Autoencoders (SAEs) (Sharkey et al., 2022; Cunningham et al., 2024; Bussmann et al., 2025) identify decoding-encoding directions driven by the reconstruction of representations after decomposing them into sparse latent factors. They typically consider a latent space larger than the embedding space and may suffer from the difficulty of overcoming noise in the representations in the form of dark matter (Engels et al., 2025). In contrast, our method is independent of feature reconstruction, unrestricted by the relative sizes between the concept and the embedding space, and unlike these prior works, it implicitly links the learned directions to how the model actually utilizes them. However, in this work, we restrict our experiments to cases where the concept space is no larger than the embedding space and leave experimental comparisons to SAEs with the opposite assumption for future work. Further discussion regarding SAEs is provided in Section A.23.
@@
Moving away from reconstructing representations, the methods of Doumanoglou et al. (2023; 2024) perform directional clustering of feature activations, a process that often unveils the decoding mechanism of interpretable, monosemantic concepts. We build on these previous approaches to estimate decoding directions, removing orthogonality and feature-standardization constraints and introducing new loss terms that preserve or improve concept interpretability while reducing the impact of distractor noise in the filter weights. Compared to the previous attempt (Doumanoglou et al., 2024) of linking directions to their utilization by the network, our Uncertainty Region Alignment approach achieves up to 22.56% relative improvement in interpretability metrics in 3 out of 4 cases. We also address the estimation of concept encoding directions, which these prior works did not address. Further details on this comparison are provided in Section A.3.
@@ 2.2 Applications of Directions
Our learned direction pairs enable a series of applications, from global model understanding and detailed spatially-aware local explanations down to model intervention, which allows for counterfactual explanations and model correction. Typical previous enablers of such applications are the Concept Bottleneck Models (CBMs) (Koh et al., 2020). Classic CBMs consist of a backbone feature extractor, a concept bottleneck predicting concept content from features, and a linear head mapping concepts to classes. What makes this architecture interpretable is the fact that the concept vector, i.e, the output of the concept bottleneck, is trained to have each dimension aligned with an interpretable concept. Recent post-hoc CBM variants (Yuksekgonul et al., 2023; Oikarinen et al., 2023) retrofit this structure onto existing black-box models by freezing the original backbone, discarding its head, inserting a concept bottleneck projection into concept space, and learning a task-specific sparse linear head over concepts. To learn concepts without annotations, these works exploit CLIP Radford et al. (2021), and further advances Benou and Raviv (2025) use visual prompting (Shtedritski et al., 2023) to make the bottleneck spatially aware, allowing localization of concepts at the patch level rather than only at the image level. Most recently (Rao et al., 2024), SAEs (Cunningham et al., 2024) have been used to build CBMs directly for CLIP.
To access the advantages of CBMs, these previous approaches require training the last linear head using the model’s training data and, sometimes, additional concept-annotated datasets or external models like CLIP. In contrast, the proposed work is non-intrusive: it doesn’t need to train new model components and does not require concept labels. It shows that one can capture many benefits of concept bottleneck models by identifying directions corresponding to concepts the model already encodes and then using these directions in applications of mechanistic interpretability (Saphra and Wiegreffe, 2024; Kästner and Crook, 2024; Bereska and Gavves, 2024).
@@ 3 Background
In our formulation, which follows the linear representation hypothesis (Elhage et al., 2022; Nanda et al., 2023), the latent factor of a concept is a scalar denoting the concept’s presence in the input, and is embedded in the latent space via multiplication with the concept’s encoding direction, also called the concept’s embedding or signal direction (Kindermans et al., 2017). For this reason, we also refer to this latent factor as the signal value. Feature activations are considered as linear combinations of signals and noisy directional components called distractors. In the proposed approach, a filter, through the inner product with an input representation, can extract the signal’s value and for this reason we term it a decoding direction. Below we provide a more formal explanation of these terms and provide details essential to understand our contributions.
@@ 3.2 Signals, Distractors, Filters, Concept-Detectors
In encoding a single concept $i$, Kindermans et al. (2017); Pahde et al. (2024) proposed a model for the data generation process of feature representations: ${\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{d}},\,{\bm{s}}_{i},{\bm{d}}\in{\mathbb{R}}^{D},\alpha_{{\bm{p}}},\beta_{{\bm{p}}}\in{\mathbb{R}}$. Here, ${{\bm{s}}}_{i}$ is the signal direction that carries the information of whether ${\bm{x}}_{{\bm{p}}}$ is part of concept $i$. The concept information lies within the signal value $\alpha_{{\bm{p}}}$. Larger $\alpha_{{\bm{p}}}$ suggests greater confidence that ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. $\bm{d}$} is the distractor direction, modeling noise, or information not related to the concept. $\beta_{{\bm{p}}}$ follows a random distribution, typically the gaussian or uniform distribution, and is independent of whether ${\bm{x}}_{{\bm{p}}}$ belongs to concept $i$. According to Kindermans et al. (2017), the value of the signal $\alpha_{{\bm{p}}}$ can be extracted using a filter ${\bm{w}}_{i}$ and the inner product: $z_{{\bm{p}},i}={\bm{w}}_{i}^{T}{\bm{x}}_{{\bm{p}}}=\alpha_{{\bm{p}}}{\bm{w}}_{i}^{T}{{\bm{s}}}_{i}+\beta_{{\bm{p}}}{\bm{w}}_{i}^{T}{\bm{d}}$, if we choose ${\bm{w}}_{i}$ : ${\bm{w}}_{i}\perp{\bm{d}}$, and ${\bm{w}}_{i}^{T}{{\bm{s}}}_{i}=1$. Let $\sigma$ denote the sigmoid function. Since stronger values of $\alpha_{{\bm{p}}}$ indicate more confidence in concept presence, when combined with a threshold $b_{i}\in{\mathbb{R}}$ that can be learned from data, this filter can be turned into a concept detector: $y_{{\bm{p}},i}=\sigma(z_{{\bm{p}},i}-b_{i})$, essentially a binary classifier that can answer the question of whether $\bm{p}$} belongs to concept $i$.
With access to the signal value, Haufe et al. (2014); Kindermans et al. (2017) offer a formula to estimate the concept’s signal direction:
@@ 3.3 Unsupervised Interpretable Direction Learning
Figure 3: The core concept of Unsupervised Interpretable Basis Extraction (UIBE Doumanoglou et al. (2023)) is to learn a set of concept detectors, which are essentially binary linear classifiers with learnable filters and biases. These detectors aim to transform feature representations to the soft-binary vector space of concepts in which the newly transformed representations are sparse. In this procedure the input to the method consists of image representations coming from an unlabeled concept dataset. Identifying the concept name behind each detector is done in a post-processing step with a procedure we refer to as Direction Labeling.
Recent research (Doumanoglou et al., 2023) introduced an unsupervised method to identify concepts from the structure of the latent space. Motivated by the directional encoding of concepts, the method partitions the latent space into linear regions, each represented by a hyperplane and a normal vector, forming clusters. Feature activations from an unlabeled concept dataset, possibly activations of images from the network’s training set, are assigned to these clusters. The method learns $\bm{W}$} and $\bm{b}$} of a feature-to-cluster membership function, a mapping to the semantic space, with ${\bm{y}}_{{\bm{p}}}=\sigma({\bm{W}}^{T}{\bm{x}}_{{\bm{p}}}-{\bm{b}})\in[0,1]^{I},{\bm{W}}\in{\mathbb{R}}^{D\times I},{\bm{b}}\in{\mathbb{R}}^{I}$, and $I$ as the cluster count. By softly assigning features to a small number of clusters, the interpretability of the clustering is improved. This is grounded in the idea that an image patch generally holds only a few semantic labels from a larger set, reflecting sparsity in the semantic space. Sparsity in the assignments is achieved using two loss terms: the first is Sparsity Loss ($\mathcal{L}^{s}$), and the second is Maximum Activation Loss ($\mathcal{L}^{ma}$), which ensures binary cluster membership: